# biblio.bib

### v2

@@ 4 Analysis of the algorithm
In each cluster $C^{(m)}_{s}$ we also consider the natural ordering of the sets composing the sequence $S^{(m)}_{s}$ of $R_{m}$-improving sets that were chosen by the leader of $C^{(m)}_{s}$ (Line 14). We can create a sequence of $R_{j^{\star}+1}$-improving sets by concatenating all these maximal sequences
$((S^{(m)}_{s})_{s\in[k_{m}]})_{j^{\star}+1\leq m\leq i},$
and at the end we put the improving graph $(A,\ell_{A})$, which has improving ratio at least $\lambda$ (while $R_{j^{\star}+1}\leq 3\lambda/4$). Now, by Lemma 4.6 we have that there exists an improving set $(A^{\prime},\ell_{A^{\prime}})$ that is fully contained in the sequence $((S^{(m)}_{s})_{s\in[k_{m}]})_{j^{\star}+1\leq m\leq i}$ and the set $(\mathcal{N}_{c_{2}\operatorname{log}^{2}n/\varepsilon^{2}}[v])$ for any $v\in A$, with $\operatorname{IR}(A^{\prime},\ell_{j^{\star}},\ell_{A^{\prime}})\geq R_{j^{\star}+1}-2\varepsilon\geq R_{j^{\star}+1}$. Since
$A^{\prime}\subseteq\mathcal{N}_{c_{2}\operatorname{log}^{2}n/\varepsilon^{2}+2r+1},$
which is, in turn, fully contained in some cluster $C^{(j^{\star})}_{s}$, the sequence $(S^{(j^{\star})}_{s})_{s\in[k_{j}]}$ was not maximal, reaching a contradiction. ∎
@@
Let $i=c_{1}\operatorname{log}n$ be the last phase, and let $\ell_{i}$ be the labeling obtained at the end of phase $i$ (Lines 18 and 23). By contradiction, suppose that there is an error in the graph, that is, a centered graph $(H,v_{H})$ of radius $2r$ that does not belong to $\mathcal{C}$. This means that $v_{H}$ can change the output labels of itself and its incident edges to get an improvement of at least $\lambda$. Let $\ell_{v}$ be that relabeling, so that we can get the improving set ($(\set{v_{H}},\ell_{v})$, with $\operatorname{IR}(\set{v_{H}},\ell_{i},\ell_{v})\geq\lambda$.
By Lemma 4.4, it must hold that $A$ is fully contained in
@@
Notice that, by Lemma 3.6, at each phase $j\leq i$, for every node $v$ there is constant probability $p>0$ that $\mathcal{N}_{t_{2}+1}[v]$ is fully contained within some cluster of the $j$-th run of MPX. Hence, there exists a large enough constant $q>0$ such that $(1-p)^{q\operatorname{log}n}\leq 1/n^{11}$. Hence, with probability $1-1/n^{11}$, $\mathcal{N}_{t_{2}+1}[v_{H}]$ was contained in some cluster $C_{j}(v_{H})$ in some phase $1\leq j\leq i$ (with $i\geq q\operatorname{log}n$—hence, $c_{1}=q$), which reaches a contradiction with Lemma 4.4.
Now, by the union bound, with probability $1-1/n^{10}$, we can exclude that all nodes are centers of centered graphs of radius $2r$ that are invalid according to $\mathcal{C}$, yielding the thesis. ∎
@@ 4.1 Technical claims and proofs
Here we prove Lemma 4.5 and Lemma 4.6. The first one basically says that, for any minimal improving set $A$ of some improving ratio $\beta$, for any node $v\in A$, we can find, within some radius-$O(\operatorname{log}n/\varepsilon)$ neighborhood of $v$, an improving set $A^{\prime}\subseteq A$ with improving ratio at least $\beta-\varepsilon$. In order to prove it, we start with a preliminary lemma.
@@
$\operatorname{Imp}(S,\ell,\ell^{\star}_{S})\leq\operatorname{Imp}(A,\ell,\ell^{\star}_{A})+\operatorname{Imp}(B,\ell,\ell^{\star}_{B})+2\cdot\Delta^{2r-1}\absolutevalue{E(A,B)}\cdot\Lambda,$
where $E(A,B)=\left\{\set{a,b}\in E\ \middle|\ a\in A,b\in B\right\}$. Where $\Lambda$ is the maximal improvement possible for a node and a labeling $\ell^{\star}_{M}$ is the best possible labeling for a set $M\subset V$.
@@
=\ = $\sum_{v\in V}\Psi_{\ell}(G_{r}(v),v)-\Psi_{\ell^{\star}_{S}}(G_{r}(v),v)$
=\ = $\sum_{v\in\mathcal{N}_{r}[S]}\Psi_{\ell}(G_{r}(v),v)-\Psi_{\ell^{\star}_{S}}(G_{r}(v),v)+\sum_{v\in V(G)\setminus\mathcal{N}_{r}[S]}\Psi_{\ell}(G_{r}(v),v)-\Psi_{\ell^{\star}_{S}}(G_{r}(v),v)$
=\ = $\sum_{v\in\mathcal{N}_{r}[S]}\Psi_{\ell}(G_{r}(v),v)-\Psi_{\ell^{\star}_{S}}(G_{r}(v),v),$
Let $\ell_{A}$ be the labeling defined as follows: On the nodes in $B$ and their incident half-edges, it is equal to $\ell^{\star}_{S}$, and on the rest of the graph it is equal to $\ell$. We similarly define $\ell_{B}$ (we just replace $B$ with $A$ in the previous definition).
We want to analyze the improvement of $(B,\ell_{A})$ and $(A,\ell_{A})$ compared to that of $(S,\ell^{\star}_{S})$. Consider any node $v$ such that its distance from $A$ is at least $r+1$. The two labelings $\ell_{A}$ and $\ell^{\star}_{S}$ will yield the same evaluation of the potential on the centered graph $(G_{r}(v),v)$. Consider any node $v$ such that its distance from $B$ is at least $r+1$. The two labelings $\ell_{B}$ and $\ell^{\star}_{S}$ will yield the same evaluation of the potential on the centered graph $(G_{r}(v),v)$.
Let
$C=\left\{v\in V\ \middle|\ \operatorname{dist}(v,A)\leq r\text{ and }\operatorname{dist}(v,B)\leq r\right\}$
be the set of nodes that are at most $r$-hops away from any node in $A$ and any node in $B$. The following holds:
$\operatorname{Imp}(B,\ell,\ell_{A})$
=\ = $\sum_{v\in\mathcal{N}_{r}[B]}\left[\Psi_{\ell_{A}}(G_{r}(v),v)-\Psi_{\ell}(G_{r}(v),v)\right]$
=\ = $\sum_{v\in\mathcal{N}_{r}[B]\setminus C}\left[\Psi_{\ell_{A}}(G_{r}(v),v)-\Psi_{\ell}(G_{r}(v),v)\right]+\sum_{v\in C\cap\mathcal{N}_{r}[B]}\left[\Psi_{\ell_{A}}(G_{r}(v),v)-\Psi_{\ell}(G_{r}(v),v)\right]$
=\ = $\sum_{v\in\mathcal{N}_{r}[B]\setminus C}\left[\Psi_{\ell^{\star}_{S}}(G_{r}(v),v)-\Psi_{\ell}(G_{r}(v),v)\right]+\sum_{v\in C\cap\mathcal{N}_{r}[B]}\left[\Psi_{\ell_{A}}(G_{r}(v),v)-\Psi_{\ell}(G_{r}(v),v)\right].$
Notice that we can get the exact same equality when replacing $B$ with $A$:
$\operatorname{Imp}(A,\ell,\ell_{A})$
=\ = $\sum_{v\in\mathcal{N}_{r}[A]\setminus C}\left[\Psi_{\ell^{\star}_{S}}(G_{r}(v),v)-\Psi_{\ell}(G_{r}(v),v)\right]+\sum_{v\in C\cap\mathcal{N}_{r}[A]}\left[\Psi_{\ell_{A}}(G_{r}(v),v)-\Psi_{\ell}(G_{r}(v),v)\right].$
$\operatorname{Imp}(A,\ell,\ell^{\star}_{A})+\operatorname{Imp}(B,\ell,\ell^{\star}_{A})$
[…]
=\ = $\sum_{v\in\mathcal{N}_{r}[A]\setminus C}\left[\Psi_{\ell^{\star}_{S}}(G_{r}(v),v)-\Psi_{\ell}(G_{r}(v),v)\right]+\sum_{v\in C\cap\mathcal{N}_{r}[A]}\left[\Psi_{\ell_{A}}(G_{r}(v),v)-\Psi_{\ell}(G_{r}(v),v)\right]$
$+$ $\sum_{v\in\mathcal{N}_{r}[B]\setminus C}\left[\Psi_{\ell^{\star}_{S}}(G_{r}(v),v)-\Psi_{\ell}(G_{r}(v),v)\right]+\sum_{v\in C\cap\mathcal{N}_{r}[B]}\left[\Psi_{\ell_{A}}(G_{r}(v),v)-\Psi_{\ell}(G_{r}(v),v)\right]$
=\ = $\sum_{v\in\mathcal{N}_{r}[S]\setminus C}\left[\Psi_{\ell^{\star}_{S}}(G_{r}(v),v)-\Psi_{\ell}(G_{r}(v),v)\right]$
$+$ $\sum_{v\in C}\left[\Psi_{\ell_{A}}(G_{r}(v),v)+\Psi_{\ell_{A}}(G_{r}(v),v)-2\Psi_{\ell}(G_{r}(v),v)\right]$
=\ = $\operatorname{Imp}(S,\ell,\ell^{\star}_{S})-\sum_{v\in C}\Psi(G_{r}(v),\ell^{\star}_{S})-\Psi_{\ell}(G_{r}(v),v)$
$+$ $\sum_{v\in C}\left[\Psi_{\ell_{A}}(G_{r}(v),v)+\Psi_{\ell_{A}}(G_{r}(v),v)-2\Psi_{\ell}(G_{r}(v),v)\right]$
=\ = $\operatorname{Imp}(S,\ell,\ell^{\star}_{S})+\sum_{v\in C}\left[\Psi_{\ell_{A}}(G_{r}(v),v)+\Psi_{\ell_{A}}(G_{r}(v),v)-\Psi_{\ell^{\star}_{S}}(G_{r}(v),v)-\Psi_{\ell}(G_{r}(v),v)\right]$
[…]
The claim follows by observing that |$\absolutevalue{C}\leq\Delta^{2r-1}\absolutevalue{E(A,B)}$ | and rearranging terms. ∎
With Lemma 4.7 in hand, we can now prove Lemma 4.5.
Proof of Lemma 4.5.
For any subset $A\subseteq S$ we again define $B=S\setminus B$. Let $x_{S}=\operatorname{IR}(S,\ell,\ell_{S})$. Starting with the inequality of Lemma 4.7 and using that $S$ is minimal, we have
$\operatorname{Imp}(A,\ell,\ell^{\star}_{A})$ $\geq\operatorname{Imp}(S,\ell,\ell^{\star}_{S})-\operatorname{Imp}(B,\ell^{\star}_{B})-2\Delta^{2r-1}\absolutevalue{E(A,B)}\cdot\Lambda$
@@
So if |$\absolutevalue{E(A,B)}\leq\frac{\varepsilon}{2\Delta^{2r-1}\Lambda}\absolutevalue{A}$ |, we have that $A$ has improving ratio at least $\beta-\varepsilon$ as desired. For the sake of readability, we set $y=2\Delta^{2r-1}\Lambda$: we need to find a set $A$ with the desired properties, such that |$\absolutevalue{E(A,B)}\leq\frac{\varepsilon}{y}\absolutevalue{A}$ |.
Fix $v\in S$ and consider the following “candidate sets” which are obtained through the ball-growing sequence
$C_{0}=\set{v},\,C_{1}=\mathcal{N}_{1}[v]\cap S,\,C_{2}=\mathcal{N}_{2}[v]\cap S,\,\ldots,\,C_{i}=\mathcal{N}_{i}[v]\cap S.$
@@
|$\absolutevalue{C_{k}}\geq\left(1+\frac{\varepsilon}{y\Delta}\right)^{k}\geq\exp\left[\frac{y\Delta\cdot M\operatorname{log}n}{\varepsilon+y\Delta}\right]\geq n^{M/2},$
where we used that $(1+x/y)^{y}\geq\exp(xy/(x+y))$ for all $x>0,y>0$: we reach a contradiction because we run out of nodes.
Then, either there is an $i\leq\left\lceil My\Delta\operatorname{log}n/\varepsilon\right\rceil$ such that $C_{i}=S$, and then we are done, or the statement holds for some $i\leq\left\lceil My\Delta\operatorname{log}n/\varepsilon\right\rceil$. ∎
@@
Now we want to prove Lemma 4.6. The key ingredient for Lemma 4.6 is the next lemma, which states the following: in a sequence of $\beta$-improving sets in $G$ w.r.t. some labeling $\ell$, for any node $v$ in any of the improving sets, there exists an improving set $(A,\ell_{A})$ in the radius-$O(\operatorname{log}^{2}n/\varepsilon^{2})$ neighborhood of $v$ (that is fully contained in the union of the improving sets) that has improving ratio at least $\beta-\varepsilon$ w.r.t. $\ell$. Now, we can take the minimal improving set $A^{\prime}\subseteq A$ and, by Lemma 4.5, we can find a set $A^{\prime\prime}\subseteq A^{\prime}$ such that $(A^{\prime\prime},\ell^{\star}_{A^{\prime\prime}})$ that has improving ratio at least $\beta-2\varepsilon$ w.r.t. $\ell$ and diameter at most $O(\operatorname{log}n/\varepsilon)$. Hence, whenever there is a sequence of $\beta$-improving sets, before any relabeling, any node in the sequence can locally find small diameter improving sets with improving ratio at least $\beta-2\varepsilon$ w.r.t. $\ell$.
@@
Let $((A_{i},\ell_{i}))_{1\leq i\leq h}$ be a sequence of $\beta$-improving sets for some $\beta>0$ in $G$ w.r.t. $\ell$. Fix any value $0<\varepsilon<\beta$, such that all $A_{i}$ have strong diameter at most $c_{3}\operatorname{log}n/\varepsilon$ ($c_{3}$ as in Line 16). Then for all $1\leq j\leq h$, for all $v\in A_{j}$ there exists a large enough constant $c_{2}>0$ and some $k\leq c_{2}\operatorname{log}^{2}n/\varepsilon^{2}$, such that there exists a set $A\subseteq\mathcal{N}_{k}[\{v\}]\cap(\bigcup_{1\leq i\leq h}A_{i})$ with $\operatorname{IR}(A,\ell,\ell^{*}_{A})\geq\beta-\varepsilon$.
Before proving Lemma 4.6, we remark that $c_{2}$ (Line 9 of the algorithm) is the hidden constant in $O(\operatorname{log}^{2}n/\varepsilon^{2})$ in the statement of Lemma 4.8.
@@
We will employ an argument that is similar to that in Lemma 4.7, but this time we will grow a ball in an auxiliary graph. We define this auxiliary graph $H=(V_{H},E_{H})$ where $V_{H}$ contains all $A_{i}$ as elements (for $i\in[h]$), and $A_{i}$, $A_{j}$ are adjacent (in $H$) if $\operatorname{dist}(A_{i},A_{j})\leq 2r$ in $G$.
@@
The reason for defining the edges like this is so that if the distance of $A_{i}$ and $A_{j}$ is at most $2r$, there are centered graphs of radius $r$ of which both the relabeling of $A_{i}$ and $A_{j}$ will affect the potential. And so conversely if $A_{i}$ and $A_{j}$ are not adjacent in $H$, then no node can witness changes in both $A_{i}$ and $A_{j}$.
@@
We again want to do a ball growing argument, where we have some exponential increase for every step. However, this time the growth is defined with respect to a weighting function. We define this weighting function $w\colon V_{H}\to\mathbb{N}$ for the nodes in $H$ by $w(x)=\absolutevalue{x}$ |, that is, to any element $x=A_{i}$ (for some $i\in[h]$), we assign a weight that is equal to the cardinality of $A_{i}$ in $G$. With standard abuse of notation, we define $w(B)=\sum_{A_{j}\in B}w(A_{j})$ for every subset $B$ of $V_{H}$.
We will define a ball growing process with an exponential growth in the weight of the balls. To argue that this process ends within a number of steps of magnitude $O(\operatorname{log}n/\varepsilon)$ (in $H$), we need that the total weight is not too large.

### v3

@@ 4 Analysis of the algorithm
In each cluster $C^{(m)}_{s}$ we also consider the natural ordering of the sets composing the sequence $(S^{(m)}_{s,t},\ell_{S^{(m)}_{s,t}})_{t\in[k_{s}^{(m)}]}$ of $R_{m}$-improving sets that were chosen by the leader of $C^{(m)}_{s}$ (Line 14), where $k_{s}^{(m)}$ is the number of minimal improving sets that have been flipped inside $C^{(m)}_{s}$ at phase $m$. We can create a sequence of $R_{j^{\star}+1}$-improving sets by concatenating all these maximal sequences
$\left(\left((S^{(m)}_{s,t},\ell_{S^{(m)}_{s,t}})_{t\in[k_{s}^{(m)}]}\right)_{s\in[h_{m}]}\right)_{j^{\star}+1\leq m\leq i}$
and at the end we put the improving set $(A,\ell_{A})$. Observe that $(A,\ell_{A})$ is such that $\operatorname{IR}(A,\ell_{i},\ell_{A})\geq R_{i}\geq R_{j^{\star}+1}$ and hence the whole sequence is an $R_{j^{\star}+1}$-improving sequence. Let us rename this sequence as $(\hat{S}_{t})_{t\in[t_{i}]}$, where $t_{i}$ is the overall length of the defined sequence.
Now, for each $v\in A$, by Lemma 4.6 we have that there exists a minimal improving set $(A^{\prime},\ell_{A^{\prime}})$ that is fully contained in the set
$\mathcal{N}_{c_{2}\operatorname{log}^{2}n/\varepsilon^{2}}[v]\cap\left(\cup_{t\in[t_{i}]}\hat{S}_{t}\right),$
and such that $\operatorname{IR}(A^{\prime},\ell_{j^{\star}},\ell_{A^{\prime}})\geq R_{j^{\star}+1}-2\varepsilon$. By Lemma 4.5, for each $v\in A^{\prime}$, there exists a minimal improving set $(A^{\prime\prime},\ell_{A^{\prime\prime}})$ such that $A^{\prime\prime}\subseteq A^{\prime}\cap\mathcal{N}_{c_{3}\operatorname{log}n/\varepsilon}[v]$ and $\operatorname{IR}(A^{\prime\prime},\ell_{j^{\star}},\ell_{A^{\prime\prime}})\geq R_{j^{\star}+1}-3\varepsilon\geq R_{j^{\star}}$. The fact that
$A^{\prime\prime}\subseteq\mathcal{N}_{c_{2}\operatorname{log}^{2}n/\varepsilon^{2}}[v]$
implies that $A^{\prime\prime}\subseteq\mathcal{N}_{c_{2}\operatorname{log}^{2}n/\varepsilon^{2}}[A]$. Since $\mathcal{N}_{c_{2}\operatorname{log}^{2}n/\varepsilon^{2}+2r+1}[A]$ is fully contained in some cluster $C^{(j^{\star})}_{s}$, so it is $\mathcal{N}_{2r+1}[A^{\prime\prime}]$. Hence, the sequence $(S^{(j^{\star})}_{s,t},\ell_{S^{(j^{\star})}_{s,t}})_{t\in[k_{s}^{(j^{\star})}]}$ was not maximal, reaching a contradiction. ∎
@@
Let $i=c_{1}\operatorname{log}n$ be the last phase, and let $\ell_{i}$ be the labeling obtained at the end of phase $i$ (Lines 18 and 23). By contradiction, suppose that there is an error in the graph, that is, a centered graph $(H,v_{H})$ of radius $2r$ that does not belong to $\mathcal{C}$. This means that $v_{H}$ can change the output labels of itself and its incident edges to get an improvement of at least $\lambda$. Let $\ell_{v}$ be that relabeling, so that we can get the improving set $(\set{v_{H}},\ell_{v})$, with $\operatorname{IR}(\set{v_{H}},\ell_{i},\ell_{v})\geq\lambda$. Let us now take the minimal version $(\set{v_{H}},\ell_{v}^{\star})$ of $(\set{v_{H}},\ell_{v})$.
Since $R_{c_{1}\operatorname{log}n}\leq\lambda$ by definition, Lemma 4.4 implies that $\set{v_{H}}$ is fully contained in
@@
Notice that, by Lemma 3.6, at each phase $j\leq i$, for every node $v$ there is probability at least $1/2$ that $\mathcal{N}_{t_{2}+1}[v]$ is fully contained within some cluster of the $j$-th run of MPX. Note that $(1/2)^{c_{1}\operatorname{log}n}=1/n^{c_{1}}$. Hence, with probability $1-1/n^{c_{1}}$, $\{v_{H}\}$ is not contained in
reaching a contradiction with Lemma 4.4.
Now, by the union bound, with probability $1-1/n^{c_{1}-1}$, there is no node that is a center of a centered graph of radius $2r$ that is invalid according to $\mathcal{C}$, yielding the thesis. ∎
@@ 4.1 Technical claims and proofs
Here we prove Lemma 4.5 and Lemma 4.6. The first one basically says that, for any minimal improving set $A$ of some improving ratio $\beta$, for any node $v\in A$, we can find, within some radius-$O(\operatorname{log}n/\varepsilon)$ neighborhood of $v$, a minimal improving set $A^{\prime}\subseteq A$ with improving ratio at least $\beta-\varepsilon$. In order to prove it, we start with a preliminary lemma.
@@
[…]
Recall that $\Lambda$ is the maximum improvement associated to $(\Pi,\Psi)$, and the labeling $\ell^{\star}_{M}$ is the best possible relabeling of $M$.
@@
Let $\ell_{A}$ be the labeling defined as follows: On the nodes in $A$ and their incident half-edges, it is equal to $\ell^{\star}_{S}$, and on the rest of the graph it is equal to $\ell$. We similarly define $\ell_{B}$ (we just replace $B$ with $A$ in the previous definition).
We want to analyze the improvements of $(B,\ell_{B})$ and $(A,\ell_{A})$ compared to that of $(S,\ell^{\star}_{S})$. Consider any node $v$ such that its distance from $A$ (resp. $B$) is at least $r+1$. The two labelings $\ell_{A}$ (resp. $\ell_{B}$) and $\ell^{\star}_{S}$ will yield the same evaluation of the potential on the centered graph $(G_{r}(v),v)$.
Let $C=\mathcal{N}_{r}[A]\cap\mathcal{N}_{r}[B]$. The following holds:
$\operatorname{Imp}(B,\ell,\ell_{B})$
@@
[…]
Notice that we can get the exact same equality when replacing $B$ with $A$:
$\operatorname{Imp}(A,\ell,\ell_{A})$
[…]
$\operatorname{Imp}(A,\ell,\ell^{\star}_{A})+\operatorname{Imp}(B,\ell,\ell^{\star}_{B})$
[…]
$+$ $\sum_{v\in\mathcal{N}_{r}[B]\setminus C}\left[\Psi_{\ell}(G_{r}(v),v)-\Psi_{\ell^{\star}_{S}}(G_{r}(v),v)\right]+\sum_{v\in C\cap\mathcal{N}_{r}[B]}\left[\Psi_{\ell}(G_{r}(v),v)-\Psi_{\ell_{B}}(G_{r}(v),v)\right]$
[…]
- $\sum_{v\in C}\left[\Psi_{\ell_{A}}(G_{r}(v),v)+\Psi_{\ell_{B}}(G_{r}(v),v)-2\Psi_{\ell}(G_{r}(v),v)\right]$
[…]
- $\sum_{v\in C}\left[\Psi_{\ell_{A}}(G_{r}(v),v)+\Psi_{\ell_{B}}(G_{r}(v),v)-2\Psi_{\ell}(G_{r}(v),v)\right]$
[…]
The claim follows by the definition of $C$ and by rearranging terms. ∎
With Lemma 4.7 in hand, we can now prove Lemma 4.5.
Proof of Lemma 4.5.
For any subset $A\subseteq S$ we again define $B=S\setminus A$. Let $x_{S}=\operatorname{IR}(S,\ell,\ell_{S}^{\star})$. Starting with the inequality of Lemma 4.7 and using that $(S,\ell^{\star}_{S})$ is minimal, we have
[…]
For the sake of readability, we set $y=2\Lambda$. Fix $v\in S$ and consider the following “candidate sets”: $C_{i}=\mathcal{N}_{i}[v]\cap S$ for $i\geq 0$. Let $|n_{i}=\absolutevalue{\mathcal{N}_{r}[C_{i}]}$, and let $k$ be the smallest integer $i\geq 2r$ such that $n_{i}\leq n_{i-2r}(1+\varepsilon/(y(\Delta^{r}+1)))$.
We now prove that such $k$ must exist and must be at most $4ry(\Delta^{r}+1)\ln n/\varepsilon+2r+1$. It holds that $n_{k-1}>n_{k-1-2r}(1+\varepsilon/(y(\Delta^{r}+1)))>\ldots>(1+\varepsilon/(y(\Delta^{r}+1)))^{\left\lfloor{(k-1)}/{2r}\right\rfloor}$. By contradiction, assume $k-1>4ry(\Delta^{r}+1)\ln n/\varepsilon+2r$. By the known inequality $(1+x/y)^{y}\geq\exp[xy/(x+y)]$ for all $x>0,y>0$, we have that
$(1+\varepsilon/(y(\Delta^{r}+1)))^{\left\lfloor(k-1)/2r\right\rfloor}$
@@
$=\$$\left(1+\frac{\varepsilon}{y(\Delta^{r}+1)}\cdot\frac{\left\lfloor(k-1)/2r\right\rfloor}{\left\lfloor(k-1)/2r\right\rfloor}\right)^{\left\lfloor(k-1)/2r\right\rfloor}$
[…]
where the last inequality follows from the fact that $\varepsilon\leq y(\Delta^{r}+1)$ (as $\varepsilon<\lambda$, and $\lambda\leq\Lambda$). Since $(k-1)/2r>2y(\Delta^{r}+1)\ln n/\varepsilon+1$, we get that
$\exp\left[\frac{\varepsilon}{2y(\Delta^{r}+1)}\cdot\left\lfloor(k-1)/2r\right\rfloor\right]\geq\exp[\ln n]=n,$
and, hence, $n_{k-1}>n$, which is a contradiction.
If $C_{k}=S$, we are done. Assume that $C_{k}$ is strictly contained in $S$ and let $A=C_{k}$ and $B=S\setminus A$. We need to upper bound the size of $\mathcal{N}_{r}[A]\cap\mathcal{N}_{r}[B]$. We first show that $\mathcal{N}_{r}[A]\cap\mathcal{N}_{r}[B]\subseteq\mathcal{N}_{r}[C_{k}]\setminus\mathcal{N}_{r}[C_{k-2r}]$, and then we proceed bounding from above the cardinality of $\mathcal{N}_{r}[C_{k}]\setminus\mathcal{N}_{r}[C_{k-2r}]$ (which is, by definition, $n_{k}-n_{k-2r}$). Let $u\in\mathcal{N}_{r}[A]\cap\mathcal{N}_{r}[B]$. Trivially, $u\in\mathcal{N}_{r}[C_{k}]=\mathcal{N}_{r}[A]$. Furthermore, since $u\in\mathcal{N}_{r}[B]=\mathcal{N}_{r}[S\setminus C_{k}]$, there exists a node $w\in S\setminus C_{k}$ such that $\operatorname{dist}(u,w)\leq r$. Since $w\notin C_{k}$ but $w\in S$, we have that $\operatorname{dist}(v,w)\geq k+1$. By the triangle inequality, we have that
$\operatorname{dist}(w,v)\leq\operatorname{dist}(u,w)+\operatorname{dist}(u,v),$
which implies that
$\operatorname{dist}(u,v)\geq\operatorname{dist}(w,v)-\operatorname{dist}(u,w)\geq k+1-r.$
[…]
@@
Now we want to prove Lemma 4.6. The key ingredient for Lemma 4.6 is the next lemma, which states the following: in a sequence of $\beta$-improving sets in $G$ w.r.t. some labeling $\ell$, for any node $v$ in any of the improving sets, there exists an improving set $(A,\ell_{A})$ in the radius-$O(\operatorname{log}^{2}n/\varepsilon^{2})$ neighborhood of $v$ (that is fully contained in the union of the improving sets) that has improving ratio at least $\beta-\varepsilon$ w.r.t. $\ell$. Now, we can take the minimal improving set $(A^{\prime},\ell_{A^{\prime}}^{\star})$ with $A^{\prime}\subseteq A$ and, by Lemma 4.5, we can find a set $A^{\prime\prime}\subseteq A^{\prime}$ such that $(A^{\prime\prime},\ell^{\star}_{A^{\prime\prime}})$ is a minimal improving set with improving ratio at least $\beta-2\varepsilon$ w.r.t. $\ell$ and diameter at most $O(\operatorname{log}n/\varepsilon)$. Hence, whenever there is a sequence of $\beta$-improving sets, before any relabeling, any node in the sequence can locally find small diameter improving sets with improving ratio at least $\beta-2\varepsilon$ w.r.t. $\ell$.
@@
Let $0<\varepsilon<\min\{\beta,\lambda\}$ be any value, for some $\beta>0$, where $\lambda$ is the minimum improvement associated to $(\Pi,\Psi)$. Let $((A_{i},\ell_{i}))_{1\leq i\leq h}$ be a sequence of $\beta$-improving sets w.r.t. $\ell$ in $G$, where each minimal improving set has diameter at most $c_{3}\operatorname{log}n/\varepsilon$ ($c_{3}$ as in Line 16). Then for all $1\leq j\leq h$, for all $v\in A_{j}$ there exists a large enough constant $c_{2}>0$ and some $k\leq c_{2}\operatorname{log}^{2}n/\varepsilon^{2}$, such that there exists a set $A\subseteq\mathcal{N}_{k}[v]\cap(\bigcup_{1\leq i\leq h}A_{i})$ with $\operatorname{IR}(A,\ell,\ell^{*}_{A})\geq\beta-\varepsilon$.
Before proving Lemma 4.8, we remark that $c_{2}$ (Line 9 of the algorithm) is given in the statement of Lemma 4.8.
@@
We will employ an argument that is similar to that of Lemma 4.7, but this time we will work on an auxiliary graph. We define the auxiliary graph $H=(V_{H},E_{H})$ where $V_{H}$ contains a node $\sigma(A_{i})$ for each $A_{i}$, $i\in[h]$, and $\{\sigma(A_{i}),\sigma(A_{j})\}\in E_{H}$ if and only if $\operatorname{dist}_{G}(A_{i},A_{j})\leq 2r$.
@@
The reason for defining the edges like this is the following: if the distance of $A_{i}$ and $A_{j}$ is at most $2r$, there are centered graphs of radius $r$ of which both the relabeling of $A_{i}$ and $A_{j}$ might affect the potential. Conversely, if $\sigma(A_{i})$ and $\sigma(A_{j})$ are not adjacent in $H$, then no node in $G$ can be affected by relabelings of both $A_{i}$ and $A_{j}$.
@@
We define a weight function $w\colon V_{H}\to\mathbb{N}$ for the nodes in $H$ by $|w(x)=\absolutevalue{\sigma^{-1}(x)}$, that is, to any element $x=\sigma(A_{i})$ (for some $i\in[h]$), we assign a weight that is equal to the cardinality of $A_{i}$. With an abuse of notation, we define $w(B)=\sum_{x\in B}w(x)$ for every subset $B$ of $V_{H}$.